The jurabib Package

Jens Berger^{∗}

                 jb@jurabib.org

Stefan Ulrich

2002/09/29 v0.52p

             CTAN: macros/latex/contrib/supported/jurabib/

Abstract

This package enables automated citation with BibTEX for legal studies and the humanities. In addition, the package provides commands for specifying editors in a commentary in a convenient way. Simplified formatting of the citation as well as the bibliography entry is also provided. It is possible to display the (short) title of a work only if an authors is cited with multiple works. Giving a full citation in the text, conforming to the bibliography entry, is supported. Several options are provided which might be of special interest for those outside legal studies—for instance, displaying multiple full citations. In addition, the format of last names and first names of authors may be changed easily. Cross references to other footnotes are possible. Language dependent handling of bibliography entries is possible by the special

language field.

Contents

1 Introduction 1

2 Prequisities 2 2.1 If you are working with a Windows-Distribution . . . . . . . . . . 3 2.2 If you are working with a Unix/Linux-Distribution . . . . . . . . . 4

3 An Example 4

4 The Tools 4 4.1 The \cite command . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4.2 The \footcite commands . . . . . . . . . . . . . . . . . . . . . . . 5 4.3 The \fullcite commands . . . . . . . . . . . . . . . . . . . . . . . 6 4.4 The \nextcite commands . . . . . . . . . . . . . . . . . . . . . . . 6 4.5 The \citefield command . . . . . . . . . . . . . . . . . . . . . . 7

5 Automations 7 5.1 Empty shortauthor-field . . . . . . . . . . . . . . . . . . . . . . . 7 5.2 Empty shorttitle-field . . . . . . . . . . . . . . . . . . . . . . . . 8 6 Options 10

∗Translated into English by Maarten Wisse.

6.1 Tweaking the Format of the Citation in the Text . . . . . . . . . . 10 6.1.1 Format of the Author and Editor . . . . . . . . . . . . . . . 10 6.1.2 Title Format . . . . . . . . . . . . . . . . . . . . . . . . . . 13 6.1.3 Separation of the editor . . . . . . . . . . . . . . . . . . . . 13 6.1.4 Behaviour in Recurring Citations . . . . . . . . . . . . . . . 14 6.1.5 Remaining Options . . . . . . . . . . . . . . . . . . . . . . . 14

6.2 Tweaking the Format of the Bibliography . . . . . . . . . . . . . . 14 6.2.1 Font Commands . . . . . . . . . . . . . . . . . . . . . . . . 14 6.2.2 Options for the Bibliography . . . . . . . . . . . . . . . . . 15 6.2.3 Further Possibilities for Customisation . . . . . . . . . . . . 16 6.2.4 Cited as . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 6.2.5 More Entry Fields and Types . . . . . . . . . . . . . . . . . 20 6.2.6 Citing Juridical Dissertations and the Like . . . . . . . . . 23

7 Remaining things … 24

8 The Configuration file jurabib.cfg 26

9 Options for Other Academic Disciplines 26

10 jurabib-specific fields 32

11 Linguistic Stuff 32

12 Across the bounderies 33 12.1 jura.cls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 12.2 bibtopic.sty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 12.3 hyperref.sty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 12.4 babel.sty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 12.5 chapterbib.sty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 12.6 bibunits.sty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 12.7 index.sty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 12.8 endnotes.sty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

13 Cooperators 35

14 Response requested … 35

1 Introduction

Up till now, support for citation according to the rules used in legal studies was only available by the environment jurabibliography from jura.cls. However, this environment does not work with BibTEX. jurabib.sty should solve this problem. The package primarily counts as an extension to jura.cls, but it is built in such a way that it works well together with all L^{A}TEX standard classes and also with the KOMA-Script-classes scrbook, scrreprt and scrartcl. Hence, it is possible to use the package together with alphanum and the book-class, for writing a doctoral dissertation.

jurabib.sty (re)defines \[foot]cite in such a way that it now contains two optional arguments, so that possible editors of a publication may be specified in the second optional argument. Apart from this, the BibTEX-styles (jurabib.bst and jureco.bst which are delivered with the package enable it to

1. automatically recognize the lastnames of the authors and cite these accordingly in short citations.

2. automatically generate short title information from article and periodical citations.

3. nevertheless explicitly specify a short form of the author’s name or title.

4. (default) give the short title only if multiple works of the author are cited. The functionality described above is exclusively carried out by the BibTEX-styles

jurabib.bst, jurunsrt.bst and jureco.bst.

2 Prequisities

After processing jurabib.ins with L^{A}TEX, the following files should be created at least:

• jurabib.sty
• jurabib.bst
• jureco.bst
• jurunsrt.bst
• dejbbib.ldf
• enjbbib.ldf
• dujbbib.ldf
• frjbbib.ldf
• spjbbib.ldf
• itjbbib.ldf
• jurabib.cfg
• jbtest.tex
• jbtestbt.tex
• jbtest.bib

In a TDS-conforming TEX-System the .sty-files and the .ldf-files should be in

/[local]texmf/tex/latex/jurabib, the .bst-files in /[local]texmf/bibtex/
bst/jurabib and the .bib-files in /[local]texmf/bibtex/bib/jurabib. May

be it’s necessary to update a “Filename Database”. If you are using teTeX, run texhash, if you are using MiKTeX, run initexmf -u on a DOS prompt or the appropriate graphical frontends. All mentioned files could also be placed in the directory in which L^{A}TEX is executed. The files jbtest.tex and jbtest.bib demonstrate the use of the jurabib-package.

jbtestbt.tex serves as a demonstration of jurabib.sty in cooperation with bibtopic.sty for generating multiple bibliographies.

One enables the package in the following way:

     \usepackage[⟨Options⟩]{jurabib}

You can use \jurabibsetup in the preamble or in the configurations file:

        \jurabibsetup{
          authorformat=smallcaps,
          commabeforerest,
          titleformat=colonsep,
          bibformat=tabular
        }

You can place several values of some keys into a pair of braces:

        \jurabibsetup{bibformat={tabular,ibidem,numbered}}

At the point where the bibliography should be placed, the following should be specified:

\bibliography{⟨BibTEX-database⟩}

followed by

     \bibliographystyle{⟨BibTEX-stylefile⟩}

Now, the package is ready to work. When you are new to BibTEX please remember that for BibTEX working properly, a sequence of one L^{A}TEX run, one BibTEX run, and two subsequent L^{A}TEX runs are required:

    latex data
   bibtex data
    latex data
    latex data

2.1 If you are working with a Windows-Distribution

You have to (!) use a large version of BibTEX (bibtex8) with a huge amount of !!! memory enabled. You can achieve that by using a command-line parameter like

--huge or --wolfgang, otherwise you will get error messages like that:

   The style file: jurabib.bst
   5017: Sorry---you've exceeded BibTeX's wizard-defined function space 3000
   (That was a fatal error)

This is problem can be solved by using one of the command line parameters

of bibtex8:
    bibtex8 --wolfgang file

2.2 If you are working with a Unix/Linux-Distribution

As a user of the teTEX- or TEXLive-Distribution you do not need bibtex8, because those BibTEX executables were compiled with enough memory enabled. But it is not senseless to use bibtex8 ! If you are using author names or titles with umlauts, bibtex8 is able to handle this by default, the normal BibTEX requires translation into {\"u} etc. At the moment only TEXlive comes with bibtex8, it is missed in teTEX. You would need to compile it yourself, you will find the sources on CTAN.

3 An Example^{1}

Suppose one wants to cite a juridical work. Then, the first thing to do is to add the work to the bibliography database:

  @BOOK{kkstrr,
    author      = {Kurt Kodal and Joachim Kr{\"a}mer},
    title       = {Stra{\ss}enrecht},
    shortauthor = {Kodal/Kr{\"a}mer},
    shorttitle  = {StrR},
    year        = 1995,
    address     = {M{\"u}nchen},
    edition     = {5.},
    pages       = {S.\,30--34, \S~24}
  }

Look at the fields shortauthor and shorttitle. These are provided by the BibTEX-styles belonging to the package. shortauthor and shorttitle are the equivalent fields available. There, the required short forms of the author and title are provided. (on this, see section 5 on page 7).

4 The Tools

4.1 The \cite command

\cite A citation is specified as usual:

   \cite[\S~12]{kkstrr}

Instead of the standard layout:

[1, § 12]

the citation now looks like:

Kodal/Krämer, § 12.

The short form of the title (StrR) would only be cited when Kodal and Krämer were cited with another work or the option titleformat=all is given. Then, the citation would look as follows:

Kodal/Krämer, StrR, § 12.

Because the command \cite is redefined by jurabib.sty, it has now two optional arguments instead of one:

1For all examples—if not explicitly specified differently—the options titleformat=commasep and commabeforerest as well as the command \cite have been used.

\cite[⟨page range⟩][⟨editor⟩]{⟨key⟩}

The sequence of optional arguments may be reversed by specifying the option

natoptargorder:

\cite[⟨editor⟩][⟨page range⟩]{⟨key⟩}

So, a possible editor will be specified as follows:

     \cite[\S~12][Bassenge]{kkstrr}

The following citation would emerge from this cite command—provided that multiple works by Kodal and Krämer are cited:

Kodal/Krämer/Bassenge, StrR, § 12.

The bibliography entry belonging to this citation would then be:

Kodal, Kurt/Krämer, Joachim: Straßenrecht, 5. Auflage, München, 1995

If it might happen that one does not want to give pageranges, margin numbers or the like, and nevertheless wants to specify an editor, then, the first optional argument should be left empty:

   \cite[][Bassenge]{kkstrr}

\citetitle The command \citetitle basically behaves like \cite, but may be used for explicitly citing by short title, independent from the author’s being cited with multiple works. Otherwise, the same goes for \citetitle as for \cite.

\cite* This asterisk-from of the \cite command cites always without title. In this case, it doesn’t matter whether the shorttitle-field has been provided in the database. The options citefull=all too, are deactivated in this case. (see sec-

! tion 6). Note, however, that hereby, ambiguous citations might occur, since the automatical specification of short title/full title is deactivated for this citation. Therefore, this command should be used with care.

\citetitlefortype This command can be used to declare a list of publication types, for those the titles should appear always:

    \citetitlefortype{article,book, ... }

\citenotitlefortype For use together with the option titleformat=all to declare a list of publication types, for those the titles shouldn’t appear:

    \citenotitlefortype{article,book, ... }

4.2 The \footcite commands

\footcite These commands are different from the \cite commands, in that they automati-

\footcite*
\footcitetitle

cally generate a footnote and place a period at the end of it. A space before the \footcite commands is ignored:

  ...  Annahme. \footcite[Rn.~357]{medicus}
  ...  Annahme.\footcite[Rn.~357]{medicus}

shows in both cases:

… Annahme.^{1} However, a combination of multiple citations should be provided as usual:

    \footnote{\cite[S.~13--34]{brox:bgb};
              \cite[S.~24]{canaris}.}

This results in:

… Annahme.^{2} Here, people should remember the period themselves, regrettably.

0.5f NEW ! If you are using some \footcite or \footnote commands subsequently, jurabib is able to put a comma between the superscripted footnote marks. This is the same behavior as known from the multiple option from the footmisc package. An example:

    ... \footcite{brox:bgb}\footcite{brox:schr}\footcite{brox:ja}

produces:

…^{1,2,3}

4.3 The \fullcite commands

\fullcite These commands generate a full citation, i.e., the complete entry from the bibliog- \footfullcite raphy is inserted here. A possibly present editor will be placed before the citation

and separated from the author by “in”.^{3} The page range will be added at the end.

4.4 The \nextcite commands

\nextciteshort With these commands you can determine with a comma-separated (without \nextcitefull whitespaces !) list of keys, how a cited work should appear from now on:

    \nextciteshort{brox:bgb,canaris, ... }

in a short or

    \nextcitefull{brox:bgb,canaris, ... }

in a long kind.

    \nextcitenotitle{brox:bgb,canaris, ... }

With that command all given works appear from now on without title. All three commands are overriding the \fullcite- and \cite* commands !

\nextcitereset With
    \nextcitereset{brox:bgb,canaris, ... }

you can switch back to the normal citation kind (specified by the chosen options).

\citeswithoutentry  The new command \citeswithoutentry is very similar to the \nextcite

commands. You are able to specify a list of works, which should not appear in the bibliography, but you can use all of the \cite commands normally. Note, that the \nextcitereset command will not work here.

1Medicus, Rn. 357.

2Brox, BGB, S. 13–34; Canaris, S. 24.

3This reflects the automatic activation of the option cofirstsep=in for this citation.

         \[foot]citep{Kraft74}  → (Kraft et al., 1937)
         \[foot]citet{Kraft74}  → Kraft et al. (1937)
         \[foot]citealt{Kraft74}  → Kraft et al. 1937
         \[foot]citealp{Kraft74}  → Kraft et al., 1937
         \[foot]citeauthor{Kraft74} → Kraft et al.
         \[foot]citeyear{Kraft74}  → 1937
         \[foot]citeyearpar{Kraft74} → (1937)

4.5 The \citefield command

With that command you have access to the contents of the fields author,

shortauthor, title, shorttitle, url, apy (Address-Publisher-Year) and year.

You have to give the field name as the first mandatory argument, the key as the second mandatory argument. Furthermore you can give a page(range) with the optional argument:

\citefield[⟨pagerange⟩]{⟨field⟩}{⟨key⟩}

With the following entry

   @BOOK{brox:bgb,
    author      = {Hans Brox},
    title       = {Allgemeiner Teil des B{\"u}rgerlichen Gesetzbuches},
    shorttitle  = {BGB~AT},
    year        = 1996,
    language    = {german},
    address     = {K{\"o}ln, Berlin, Bonn, M{\"u}nchen},
    edition     = 20
   }

we get with a \citefield{title}{brox:bgb}: Allgemeiner Teil des Bürgerlichen

Gesetzbuches. \citefield[\S~23]{shorttitle}{brox:bgb} gives: BGB AT, § 23.

This command is especially for non-lawyers. It is independent of all jurabib-specific automatisms, e.g. it is ignored by the ibidem-options. If you are using the hyperref-package, a link to the bibliography entry will be created.

jurabib is able to emulate some basic citation commands of the natbib package:

5 Automations

5.1 Empty shortauthor-field

The package is now capable of automatically figuring out the lastname of the author, in case of empty or failing shortauthor-field. Up till three authors are typed out, separated by a slash. In case of more than three authors, the name of the first author is typed out, along with “et al.”. So, for instance, we leave the shortauthor-field out in our example entry and add a third author:

  @BOOK{kkstrr,
    author      = {Kurt Kodal and Joachim Kr{\"a}mer
                   and Hans Mustermann},
    title       = {Stra{\ss}enrecht},

    shorttitle  = {StrR},
    year        = 1995,
    address     = {M{\"u}nchen},
    edition     = {5.},
    pages       = {S.\,30--34, \S~24}
  }

Then, the citation looks as follows:

Kodal/Krämer/Mustermann, StrR, § 12.

Separation marks may be modified in the following way.

   \renewcommand{\jbbtasep}{ and } % bta  = between two authors sep
   \renewcommand{\jbbfsasep}{, }   % bfsa = between first and second author sep
   \renewcommand{\jbbstasep}{ and }% bsta = between second and third author sep

shows, for instance, instead of the example shown above:

Kodal, Krämer and Mustermann, StrR, § 12.

Separation marks in the bibliography remain untouched by this modification. Those can be modified by the following commands:

   \renewcommand{\bibbtasep}{ and } % bta  = between two authors sep
   \renewcommand{\bibbfsasep}{, }   % bfsa = between first and second author sep
   \renewcommand{\bibbstasep}{ and }% bsta = between second and third author sep
and for the editors:
   \renewcommand{\bibbtesep}{ and } % bte  = between two editors sep
   \renewcommand{\bibbfsesep}{, }   % bfse = between first and second editor sep
   \renewcommand{\bibbstesep}{ and }% bste = between second and third editor sep

If we add a fourth author (we save work by skipping the code), then the citation looks like:

Kodal et al., StrR, § 12.

5.2 Empty shorttitle-field

If the shorttitle-field is empty or nonexistent, the following happens: The full title specified in the title-field will be used^{4}, if (locally) the command \[foot]citetitle or (global) the option citefull=all is being used or multiple works of the author have been cited.

In case one wants author’s name only for a particular work—regardless what kind of option have been activated—then, the command \[foot]cite* should be used. This may lead to ambiguous citations, of which you will get a warning in the .log-file.

  @BOOK{kkstrr,
    author      = {Kurt Kodal and Joachim Kr{\{\"a}}mer
                   and Hans Mustermann},
    title       = {Stra{\ss}enrecht},
    year        = 1995,
    address     = {M{\"u}nchen},
    edition     = {5.},
    pages       = {S.\,30--34, \S~24}
  }

4Exceptions are @ARTICLE- and @PERIODICAL entries.

Result:

Kodal/Krämer/Mustermann, Straßenrecht, § 12.

An alleviation should also be the fact that, in case of Article and Periodical citations, jurabib now automatically cites (⟨journal⟩ ⟨year⟩). Therefore, in these cases, specification of shorttitle is no longer required—if no special requirements are present. In the following example, neither shortauthor nor shorttitle has been specified:

   @ARTICLE{brox:ja,
     author      = {Hans Brox},
     title       = {Die Anfechtung bei der Stellvertretung},
     journal     = {JA},
     year        = 1980,
     pages       = {S.\,449ff},
     edition     = {20.},
     address     = {M{\"u}nchen}
   }

Nevertheless, \citetitle{brox:ja} results in:

Brox, JA 1980.

If you want something different, you can use the short*-fields:

   @ARTICLE{brox:ja,
     author      = {Hans Brox},
     title       = {Die Anfechtung bei der Stellvertretung},
     journal     = {JA},
     shortauthor = {Hans Brox},
     shorttitle  = {JA},
     year        = 1980,
     pages       = {S.\,449ff},
     edition     = {20.},
     address     = {M{\"u}nchen}
   }

and you get:

Hans Brox, JA.

author = You should notice that the dynamic generations of the juridical short titles is possible only, when the contents of the author-fields in the .bib-file are absolutely identical ! Suppose Kodal and Krämer have published a second work, then, the entries in the database should look as follows:

   @BOOK{kkstrr,
      author     = {Kurt Kodal and Joachim Kr{\"a}mer},
      title      = {Stra{\ss}enrecht}
   }
  @BOOK{kkanything,
      author     = {Kurt Kodal and Joachim Kr{\"a}mer},
      title      = {anything}
   }

6 Options

The following default behavior of the package has been implemented: The author and editor in the citation are shown in normal font, in the bibliography, however, in bold font. The editor is put after the author, separated by a slash. This default behavior can be modified by the options of the package.

6.1 Tweaking the Format of the Citation in the Text

6.1.1 Format of the Author and Editor

authorformat=smallcaps Puts author and editor in small caps: Kodal/Krämer/

Bassenge, StrR, § 12 authorformat=italic Puts author and editor in italic font: Kodal/Krämer/

Bassenge, StrR, § 12 authorformat=dynamic The font of the author depends on whether an editor is

present or not. If not, the authors retain their font as specified in the command \jbactualauthorfont as \textit (default): Kodal/Krämer, StrR, § 12. If an editor has been specified, the editor is formatted according to the setting of \jbactualauthorfont, and the authors are formatted according

to \jbauthorfontifcoauthor ( \normalfont): Kodal/Krämer/Bassenge,

StrR, § 12.

authorformat=citationreversed Using this option together with options citefull=first,

ibidem=name and ibidem=name&title firstnames appear before last names: Hans Brox: BGB AT, S. 23. instead of Brox, Hans: BGB AT, S. 23.

authorformat=allreversed By this option, the behavior mentioned in the pre-

vious option works also in the bibliography. authorformat=firstnotreversed Sometimes, strange requirements are set on

authors. Therefore, this option exists, which puts all authors in ⟨Firstname⟩ ⟨Lastname⟩, apart from the first author: Kodal, Kurt/Joachim Krämer/ Hans Mustermann.

authorformat=reducedifibidem If this option is active together with ibidem=name

only the last name of the author is shown in recurring citations. authorformat=and Instead of the default slashes the authors will be separated

by “,” and “, and”. authorformat=year The year will appear after the author name. \jbyearaftertitle

puts the year after the title. Formatting is possible by using \jbcitationyearformat:

  \renewcommand{\jbcitationyearformat}[1]{(#1)}

authorformat=indexed All authors (independently) are indexed. You have to use the makeidx-package correctly:

    \usepackage{jurabib} % load before makeidx.sty!
    \usepackage{makeidx}
    \makeindex
    \begin{document}
     ...
    \printindex
    \end{document}

This options works for \nobibliography too (since 0.51). If you want to have not cited authors (e.g. by using \nocite) indexed, you can use \jbindexbib in your preamble.

0.52 NEW ! To emphasize the authors in the index, you can determine the font of the

indexed authors via \jbauthorindexfont:
    \renewcommand{\jbauthorindexfont}{\textit}% or \textsf,
                                              %    \textsc,
                                              %    \textbf

If you are using a non-standard .ist file (makeindex style file), it is possible that you have to use \jbmakeindexactual to use the correct ‘actual’ operator. The default is @. If your .ist file contains

     actual '='

you should use:

     \renewcommand{\jbmakeindexactual}{\=}%

I don’t know why the backslash is necessary. authorformat=abbrv Since version 0.5 the BibTEX-styles jurplain.bst and

jurabbrv.bst are obsolete, the abbreviated form can be determined by using this option.

coauthorformat=italic Puts the editor in italics: Kodal/Krämer/Bassenge. coauthorformat=normal Puts the editor upright: Kodal/Krämer/Bassenge. round Round brackets around (non-footnote-) citation: (Kodal/Krämer). square Square brackets around (non-footnote-) citation: [Kodal/Krämer]. superscriptedition=all Places a superscripted edition number in the citation.

Some examples: Kodal/Krämer^{3}, § 12. Kodal/Krämer, Stra”senrecht^{3}, § 12. Kodal/Krämer^{3}–Bassenge, § 12. Kodal/Krämer–Bassenge, Straßenrecht^{3}, § 12. Bassenge in Kodal/Krämer^{3}, § 12.

superscriptedition=commented Same like above, but only for @COMMENTED.
superscriptedition=switch With the new field ssedition you can specify the

appearence of the superscripted edition number explicitly:

   @COMMENTED{soergel,
      author      = {Hans Theodor Soergel},
      title       = {Kommentar zum B{\"u}rgerlichen Gesetzbuch},
      address     = {Stuttgart, Berlin, K{\"o}ln, Mainz},
      year        = 1987,
      edition     = 12,
      ssedition   = 1, <=== !
   }

0.51h NEW ! superscriptedition=kerning This option replaces the command \jbsseditionkerned

and does some kerning:

Kodal/Krämer^{3}, § 12. Kodal/Krämer, Stras̈enrecht^{3}, § 12. Kodal/Krämer^{3}–Bassenge, § 12. Kodal/Krämer–Bassenge, Stras̈enrecht^{3}, § 12. Bassenge in Kodal/Krämer^{3}, § 12.

0.51j CHANGED ! superscriptedition=bib/address superscripted edition numbers in the bibli-

ography (before address).

0.51j NEW ! superscriptedition=year superscripted edition numbers in the bibliography

(before year).

0.51h NEW ! superscriptedition=multiple This option enables the output of a superscripted edition number for the case that different editions of the same work were cited. The output of the shorttitle is suppressed.

If a coauthor was given via the second optional argument, the superscripted edition number appears after the coauthor, as long as you are using the default or one of the colastsep options:

Kodal/Krämer–Bassenge^{3}, § 12.

If you like to have the superscripted number immediately after the author, please use the following command in the preamble of your document:

\jbsuperscripteditionafterauthor

Should result in:

Kodal/Krämer^{3}–Bassenge, § 12.

biblikecite Bibliography will be automatically formatted like the citations (as

far as possible !). edby (Only for @INCOLLECTION!) The sequence “Lipcoll, David J. (ed.)” will be

changed to “edited by Lipcoll, David J.”

For redefinitions please use:

   \AddTo\bibsgerman{%
       \def\edbyname{ed. by}%
    }

endnote The note field appears at the end of the bibliographic entry, and if you are using \jbdotafterbibentry, after the closing period.

\jbauthorfont If these options do not generate the desired result, one can realize many things \jbcoauthorfont by redefining several commands.

The formats of the authors and editors are directed by the commands \jbauthorfont and \jbcoauthorfont and may be modified in the following way (these examples presuppose the default values):

\renewcommand{\jbauthorfont}{\textit}
\renewcommand{\jbcoauthorfont}{\textsl}

\jbactualauthorfont Analogically, the following commands exist only if option authorformat=dynamic

\jbauthorfontifcoauthor is active:
  \renewcommand{\jbactualauthorfont}{\textsc}
  \renewcommand{\jbauthorfontifcoauthor}{\textsl}

These redefinitions should be placed in the preamble of the TEX-file, i.e. before \begin{document}. One should notice that \text commands should be used, for example, \textit, \textbf etc.), and not those starting with series, family or ending with shape (declaration form, for instance, \bfseries, \slshape,

\sffamily)!

6.1.2 Title Format

titleformat=italic Puts title in italics: Kodal/Krämer/Bassenge, StrR, § 12. titleformat=all Provides short titles in all cases, regardless whether an author

has been cited with multiple works. titleformat=colonsep Separates author and title by a colon (only if a title is

displayed): Kodal/Krämer/Bassenge: StrR, § 12. titleformat=commasep Separates author and title by a comma: Kodal/Krämer/

Bassenge, StrR, § 12. titleformat=noreplace You can disable globally the default replacement of the

     missing shorttitle by title. See also the \cite*- and \nextcitenotitle

commands.

\jbtitlefont For changing the format of the short title, the command \jbtitlefont is available, which may be tweaked analogically:

   \renewcommand{\jbtitlefont}{\textit}
\jbhowsepbeforetitle  For options titleformat=commasep and titleformat=colonsep is specific

formatting of the separation marks possible by:

   \renewcommand{\jbhowsepbeforetitle}{; } .

One of both options mentioned above should be active in this case.

6.1.3 Separation of the editor

As explained above, the format of the editor after the author(s) and the separation by a slash have been set by default. These may be modified by options.

colastsep=divis This option changes the slash as separation mark to an (en)-

dash: Kodal/Krämer–Bassenge StrR, § 12. cofirstsep=in The editor appears first in the citation, followed by “ in ” : Bas-

senge in Kodal/Krämer StrR, § 12. cofirstsep=comma This option works like cofirstsep=in, but “ in ” is now

replaced by a comma: Bassenge, Kodal/Krämer StrR, § 12.

6.1.4 Behaviour in Recurring Citations

ibidem or ibidem=strict If an author is cited several times after each other, it may be useful to replace the short citation by the shorthand “ibid.” The use of this shorthand is not accepted by everyone because it does not improve readability of the text. In this respect, jurabib has been configured in such a way that ‘ibid’ will be displayed only if the recurrence immediately follows the previous citation, and the recurring citation is not the first on the actual page (see the examples). By contrast, in the humanities the use of “ibid.” is often compulsory: it can be an error to repeat a citation in full. Because of this fact, ibidem=strict is the default for the humanities.

ibidem=nostrict If one wants to suppress the settings mentioned above, one could use this option, which allows the shorthand be placed as the first on a page. One should be careful with this option. It is useful only if one uses basically one author in the document, so that it is clear to whom one refers.

ibidem=strictdoublepage Granted, this option is of an academic nature, because it quite carefully sorts out whether an first citation or short citation occurs on the facing page of the recurring citation.(Many thanks to Stefan Ulrich) If so, ibidem is allowed as first citation on the page. If not, short citation is used for recurring citations. Of course, this is useful only when two-side printing is concerned, i.e. if the book-class is used or the class option twoside. Compare the following survey:

\footcite..       ibidem/ibidem=strict   ibidem=strictdoublepage   ibidem=nostrict

..[\S~12]{erm}   ^{1}Erman, § 12.  ^{1}Erman, § 12.  ^{1}Erman, § 12.
..[\S~12]{erm}   ^{2}ibid.  ^{2}ibid.  ^{2}ibid.
..[\S~20]{erm}   ^{3}ibid., § 20.  ^{3}ibid., § 20.  ^{3}ibid., § 20.
..{mueko}        ^{4}MüKo.  ^{4}MüKo.  ^{4}MüKo.
..[\S~12]{erm}   ^{5}Erman, § 12.  ^{5}Erman, § 12.  ^{5}Erman, § 12.
..[\S~12]{erm}   ^{6}ibid.  ^{6}ibid.  ^{6}ibid.

Pagebreak from odd (right) to even (left)

..[\S~12]{erm}   ^{7}Erman, § 12.  ^{7}Erman, § 12. ⟵ !!!  ^{7}ibid. ⟵ !!!
..[\S~12]{erm}   ^{8}ibid.  ^{8}ibid.  ^{8}ibid.

Pagebreak from even (left) to odd (right)

..[\S~12]{erm}   ^{9}Erman, § 12.  ^{9}ibid. ⟵ !!!  ^{9}ibid. ⟵ !!!
..[\S~12]{erm}   ^{10}ibid.  ^{10}ibid.  ^{10}ibid.

\noibidem Disables the ibidem-mechanism for the next (and only for the next !) citation.

6.1.5 Remaining Options

commabeforerest If active, a comma will be placed before page ranges, margin numbers or the like: Kodal/Krämer/Bassenge: StrR, § 12.

6.2 Tweaking the Format of the Bibliography

6.2.1 Font Commands

\biblnfont The possibilities for formatting are limited to the modification of font formats

\bibelnfont
\bibfnfont
\bibefnfont
\bibtfont
\bibbtfont
\bibjtfont
\bibapifont
\bibsnfont

of particular parts of entries in the bibliography. For this purpose, the following commands are available. \biblnfont, for formatting the last name of the author and \bibelnfont for the last name of the editor. \bibfnfont, for formatting the first name of author and \bibefnfont for the first name of editor. \bibtfont, for modification of the title of books etc. \bibbtfont, for formatting the title of collections of essays. \bibjtfont, for formatting the title of the journal in article entries. With \bibsnfont you can determine the appearence of the series name.

Additionally available are \bibapifont, for formatting the title of an article or essay in a collection. This command is active for the entry types @ARTICLE, @PERIODICAL and @INCOLLECTION. The default format reflects the following definitions:

   \renewcommand{\biblnfont}{\bfseries}
   \renewcommand{\bibfnfont}{\bfseries}
   \renewcommand{\bibtfont}{}
   \renewcommand{\bibbtfont}{}
   \renewcommand{\bibjtfont}{}
   \renewcommand{\bibapifont}{}

Modification is possible analogically to the examples shown above. You should notice here, too, that commands should be used which start with \text (Font switch commands with arguments, for instance, \textit, \textbf etc.), and not those starting with series, family or ending with shape (declaration form, for

instance, \bfseries, \slshape, \sffamily)!

6.2.2 Options for the Bibliography

bibformat=nohang Hereby, the default indent of the second and following lines in a bibliography entry will be suppressed. If one wants to set the indent to a certain length, one should put the following in the preamble of one’s document:

        \setlength{\jbbibhang}{1.5em}

An indent of 2.5 em is the default. bibformat=tabular In this case, the bibliography will be displayed in two-column

tabular form. The authors appear in the left column, and the remainder of the entry in the right column. The width of the columns may be customized by the following commands (the values indicated are the defaults):

         \renewcommand{\bibleftcolumn}{\textwidth/3}
         \renewcommand{\bibrightcolumn}{\textwidth-\bibleftcolumn-1cm}

Modification of the alignment within the columns is also possible by redefining the following (defaults are displayed):

         \renewcommand{\bibleftcolumnadjust}{\raggedright}
         \renewcommand{\bibrightcolumnadjust}{\raggedright}

For better hyphenation, use of the package ragged2e.sty is highly recommended:

        \usepackage{ragged2e}

Loading the package ragged2e is sufficient. Redefinition of the necessary commands is handled automatically.

bibformat=numbered This option results in a numbered bibliography. The format of the number can be determined by redefining \bibnumberformat:

  \renewcommand{\bibnumberformat}[1]{(#1)}

bibformat=ibidem Replaces recurring authors name(s) by a dash (or whatever you want), if multiple works of the author appearing in the bibliography. jurabib is built in such a way, that the replacement is suppressed, when a recurring entry lies on top of a page. Note that it may be necessary to do several (up to four or more !) L^{A}TEX-runs to make this mechanism work.

Modifications can be done by using (only an example !):

  \renewcommand{\bibauthormultiple}{The same}

bibformat=ibidemalt An alternative format of the bibliography, especially for

German law students. (suggested by Tilman.Finke@web.de). bibformat=compress The bibliography will be printed more compact, i. e. the

vertical space between the items will be reduced. bibformat=raggedright The bibliography will be printed with right ragged mar-

gin. The use is recommended especially when using bibformat=tabular too or when you are using a small textwidth.

annote The content of the annote field will be printed (only for the bibliogra-

phy !). It is possible – similar to natbib – to leave the annote field empty

0.51s NEW ! and let jurabib use an annote file instead. This file will be used if it is named like the database entry key with .tex extension. Nothing will be printed out, if annote field is empty and no annote file exists.

\bibAnnotePath With \bibAnnotePath you can specify a path to annote files. The syntax

is the same as for \graphicspath: \bibAnnotePath{{annotes/}} uses the

annote files from subdirectory annotes of the current directory.

6.2.3 Further Possibilities for Customisation

\bibbtsep Because of usually different opinions about what citations should look like, the \bibjtsep commands \bibbtsep and \bibjtsep are available. They function as “booktitle

separation” and “journaltitle separation”.

Brinkmann, Franz Josef: Der Zugang der Willenserklärungen, Münsterische Beiträge zur Rechtswissenschaft, Bd. 3 Berlin, 1984

If one, for instance, wants “in ” preceding book titles or journal titles, then, the following redefenitions are required:

\renewcommand{\bibbtsep}{in }
\renewcommand{\bibjtsep}{in }

After redefinition, the following comes out:

Brinkmann, Franz Josef: Der Zugang der Willenserklärungen, in Münsterische Beiträge zur Rechtswissenschaft, Bd. 3 Berlin, 1984 \bibansep The separation marks between authors, titles and between address and month/

\bibatsep
\bibbdsep

year are a matter of discussion. This problem is countered by the commands \bibansep (after name separation), \bibatsep (after title separation) and \bibbdsep (before date separation).

If one, for instance, wants no colon after the author’s name, a period after the title and no comma between address and year, one could accomplish this by the following redefinitions:

  \renewcommand{\bibansep}{}
  \renewcommand{\bibatsep}{.}
  \renewcommand{\bibbdsep}{}

The result is as follows:

Brinkmann, Franz Josef Der Zugang der Willenserklärungen. Münsterische Beiträge zur Rechtswissenschaft, Bd. 3 Berlin 1984

6.2.4 Cited as …

howcited=normal The howcited-options put (under certain conditions) a commentary behind se-

howcited=multiple
howcited=compare
howcited=all

lected entries, which indicates how the work has been cited in the text. The commentary changes dynamically, just as we know that from the citation itself, i.e. the form in the bibliography always reflects the final form of the citation in the text. In case of article and periodical citations, the default is that no indication of the way of citation is given—this is the default for all howcited-options—except of howcited=all—because the way of citation is in these cases always author’s name and journal title.^{5}

If you want to have the howcited-remark for articles and periodicals too:

      \makeatletter
      \jb@allow@howcited@art@periodtrue
      \makeatother

These options can be enabled in the following way:

howcited=  howcited=normal The option howcited=normal displays the remark “(cited:

⟨author⟩)”, if in the .bib-file the field howcited has been specified.^{6} Two possibilities exist for utilizing this field. The field functions as a switch when putting 1 in the field. This results in displaying the original citation from the text in the bibliography. An example:

          @BOOK{enne:nipp,
            author      = {Ludwig Enneccerus and Hans Carl Nipperdey},
            title       = {Allgemeiner Teil des B{\"u}rgerlichen Rechts},
            year        = 1960,
            volume      = 1,
            address     = {T{\"u}bingen},
            edition     = 15,
            howcited    = 1
          }

5Special thanks to Christian Meyn for this suggestion.

6By this option, one can avoid unwanted automation of howcited remarks which may occur when using the option howcited=compare. Thus, it is possible to specify for each work whether it should have a howcited remark and if it must have, what it should look like.

Displays the following (use of \[foot]cite presupposed):

Enneccerus, Ludwig/Nipperdey, Hans Carl: Allgemeiner Teil des Bürgerlichen Rechts. Bd. 1, 15. Auflage, Tübingen 1960 (cited: Eneccerus/Nipperdey)

In case you want to let the howcited remark differ from the original citation, then, you should simply put in the field what you want to be displayed. An example:

    @BOOK{enne:nipp,
      author      = {Ludwig Enneccerus and Hans Carl Nipperdey},
      title       = {Allgemeiner Teil des B{\"u}rgerlichen Rechts},
      year        = 1960,
      volume      = 1,
      volumetitle = {zweiter Halbband},
      address     = {T{\"u}bingen},
      edition     = {15.},
      howcited    = {Enneccerus/Nipperdey, B{\"u}rgerliches Recht}
    }

Shows:

Enneccerus, Ludwig/Nipperdey, Hans Carl: Allgemeiner Teil des Bürgerlichen Rechts. Bd. 1, zweiter Halbband, 15. Auflage, Tübingen 1960 (cited: Enneccerus/Nipperdey, Bürgerliches Recht)

In order to ensure consequent layout when using options which affect font format of author’s names, one could insert fontcommands in the field.

howcited=compare howcited=compare Displays the additional “(cited: ⟨author⟩)” only, if an entry contains the field shorttitle, and the information in this field differs from that specified in title. The field howcited in the .bib-file will now be ignored. Decisive for displaying the remark is now only the difference between shorttitle and title ! That is the case in the following example, cited

with \[foot]citetitle:
    @BOOK{kkstrr,
      author      = {Kurt Kodal and Joachim Kr{\"a}mer},
      title       = {Stra{\ss}enrecht},
      shorttitle  = {StrR},
      year        = 1995,
      address     = {M{\"u}nchen},
      edition     = {5.},
      pages       = {S.\,30--34, \S~24}
    }

Kodal, K./Krämer, J.: Straßenrecht, 5. Auflage München, 1995 (cited: Kodal/Krämer, StrR)

howcited=multiple howcited=multiple This option places “(cited: ⟨author⟩)”, if more than one work of an author is cited. There is an exception for commentaries by default (the remark will be displayed always), but this could be changed (\jb@@arg=1, if more than one work of the same author is cited):

          \makeatletter
          \renewcommand{\jb@make@howcited@multiple}{%
             \jb@suppress@dot@for@howcitedtrue
             \ifthenelse{\equal{\jb@@arg}{1}}{%
               \jb@make@howcited
               \jb@make@comment@howcited
               \jb@make@artperiod@howcited
             }{%
               \let\bibhowcited\@empty
               \let\bibcommenthowcited\@empty
               \let\bibartperiodhowcited\@empty
             }%
          }%
          \makeatother

howcited=all  howcited=all The howcited-remark appears for all entries.

The default for the remark is “(cited: ⟨author⟩)”. Customization is possible by the following commands:

   \newcommand*{\bibhowcitedprefix}{-- as }
   \newcommand*{\bibhowcitedsuffix}{ cited.}

Notice the spaces used. This results in:

Kodal, K./Krämer, J.: Straßenrecht, 5. Auflage München, 1995 —as Kodal/ Krämer, StrR cited.

If you want to put something in the bibliography, but it has not been cited in the text, you can use command \nocite:

   \nocite{kkstrr}

Or, in order to put all works which are not cited in the bibliography:

   \nocite{*}

Then, the commentary contains the actual meaning of the command \bibnotcited. This command is empty by default. It might easily be modified by:

   \renewcommand{\bibnotcited}{(not cited)}

Shows the following:

Kodal, K./Krämer, J.: Straßenrecht, 5. Auflage München, 1995 (not cited)

Of course, for this to happen, one of the two option howcited=normal or

howcited=compare should be active.

0.51 NEW ! Now all useful combinations of howcited options are possible.

Remark: using these options without using the option citefull=all or for some citations the command \[foot]citetitle, seems not useful.

6.2.5 More Entry Fields and Types

url = Although citing the World Wide Web is not widely practised in law studies, I

nevertheless added an extra field url.

urldate = A field urldate is available, which enables one to specify the date on which one visited the link which has been specified in url. The default for this command is “visited on ” and may be customized by redefining:

   \AddTo\bibsenglish{\renewcommand*{\urldatecomment}{accessed on }} .

The separation between URL and \urldatecomment is configurable and represented by \bibbudcsep. It is defined as ”‘ – ”’ by default.

\biburlprefix One can customize the format of the url-field in two ways. On the one hand,

\biburlsuffix
\biburlfont

by modifying \biburlprefix, which inserts the prefix “URL:” before the link. On the other hand, by modifying the command \biburlfont, which specifies the font format of the link. Defaults are:

   \renewcommand*{\biburlprefix}{\jblangle{}URL:}
   \renewcommand*{\biburlsuffix}{\jbrangle{}}

Customisation is analog to the other commands by using \renewcommand. In order to properly break URLs and properly display characters like ~ and _, jurabib is loading the url-package. The \biburlfont command was changed in version

0.51 CHANGED ! 0.51. You can modify the url font with the following syntax (only these four values are possible !)

   \biburlfont{tt}   % typewriter
   \biburlfont{rm}   % roman
   \biburlfont{sf}   % sans serif
   \biburlfont{same} % same as text

@WWW New entry type for URL’s. Required is only url, optional are urldate, author,

and title.

@PERIODICAL After a suggestion by Andreas Stefanski, I’ve added a new entry type @PERIODICAL for periodicals which are not cited by year, but by volume number. This entry type satisfies the requirement to put the year between square brackets. Additionally, the specification of the volume is possible:

  @PERIODICAL{oellers,
    author      = {Bernd Oellers},
    title       = {Doppelwirkung im Recht?},
    journal     = {AcP},
    year        = 1969,
    volume      = 169,
    pages       = {S.\,67ff}
  }

This shows the following:

Oellers, Bernd: Doppelwirkung im Recht? AcP 169 [1969], S. 67ff

\bibpldelim Changing the format of the brackets is possible by redefining the commands \bibprdelim \bibpldelim (periodical left delimiter) and \bibprdelim (periodical right de-

limiter):

   \renewcommand{\bibpldelim}{(}
   \renewcommand{\bibprdelim}{)}

Oellers, Bernd: Doppelwirkung im Recht? AcP 169 (1969), S. 67ff

@COMMENTED By definition of the entry type @COMMENTED it is possible to cite commentaries as such. In connection with the option howcited=normal (cited as ⟨author⟩/editor) or (cited as editor in ⟨author⟩) appears at the end of the bibliography entry.

Münchener Kommentar: Kommentar zum Bürgerlichen Gesetzbuch, Bd. 2, – Schuldrecht – Allgemeiner Teil, 3. Auflage, München, 1994, §§ 241–432 (cited: MüKo/Bearbeiter)

updated = This new field will be recognized while using @COMMENTED to give the date of 0.51e NEW ! last update.

M”unchener Kommentar: Kommentar zum B”urgerlichen Gesetzbuch, Bd. 2, – Schuldrecht – Allgemeiner Teil, 3rd edition, M”unchen, last update: May 1994

updated does not overwrite the year field ! The separation from the address/ publisher/year block is done by the \updatesep macro, which is defined as comma by default. In front of the content of the updated field appears “last update”. This is configurable via \updatename.

   \AddTo\bibsenglish{%
        \def\updatesep{.}
        \def\updatename{updated:}
   }

If one doesn’t use this entry type, although the entry is a commentary, (cited as ⟨author⟩ appears, which is incorrect, because the actual citation looks different. Those who do not use the option howcited=normal cannot avoid using the entry

type @COMMENTED.

volumetitle By the field volumetitle it is possible to specify a volume title which appears after the volume number. This field is available for the entry types @COMMENTED,

@BOOK, @INBOOK and @INCOLLECTION.

titleaddon = This field can be used to place a commentary, a note, some remarks about translators, coauthors etc. after the title.

  @COMMENTED{mueko,
    [...]
    title       = {Kommentar zum B{\"u}rgerlichen Gesetzbuch},
    titleaddon  = {Unter Mitarbeit von Hans Mustermann},
    [...]
  }

shows us:

Münchener Kommentar: Kommentar zum Bürgerlichen Gesetzbuch, Unter Mitarbeit von Hans Mustermann, Bd. 2, – Schuldrecht – Allgemeiner Teil,

3. Auflage, München, 1994, §§ 241–432

booktitleaddon =  The same as titleaddon for booktitles in @INCOLLECTION’s.

editortype = If you want to place something other than “(eds.)” after a person, which isn’t really an editor, you can use the field editortype:

   @COMMENTED{palandt,

     editor      = {Otto Palandt},
     editortype  = {Begr.},
     title       = {B{\"u}rgerliches Gesetzbuch mit Einf{\"u}hrungsgesetz [...]},
     [...]
   }

Palandt, Otto (Begr.): Bürgerliches Gesetzbuch mit Einführungsgesetz [...], 59. Auflage, München, 2000

This works for @INCOLLECTION too.

sortkey = It seems to be required sometimes to determine the sorting of some works different from the normal sorting algorithm. This problem can be solved with the sortkey field, which can be used to sort the work with highest priority – contrary to the standard key field, which is mostly a fallback if author and editor are missing.

annote = Some people wrote me they would need the annote field to give a short abstract or something similar at the end of the bibliographic entry. This is working now. You are able to switch this feature on and off by using the annote option in the preamble or in your local jurabib.cfg file.

   @COMMENTED{palandt,
     editor      = {Otto Palandt},
     title       = {B{\"u}rgerliches Gesetzbuch mit Einf{\"u}hrungsgesetz [...]},
     annote      = {Some people wrote me they ...},
     [...]
   }

Palandt, Otto: Bürgerliches Gesetzbuch mit Einführungsgesetz [...], 59. Auflage, München, 2000

Some people wrote me they would need the annote field to give a short abstract or something similar at the end of the bibliographic entry. This is working now. You are able to switch this feature on and off by using the annote option in the preamble or in your local jurabib.cfg file.

The content of the field is printed out in \small by default. If you want change that:

   \renewcommand*{\jbannoteformat}[1]{{\footnotesize\begin{quote}#1\end{quote}}}

Palandt, Otto: Bürgerliches Gesetzbuch mit Einführungsgesetz [...], 59. Auflage, München, 2000

Some people wrote me they would need the annote field to give a short abstract or something similar at the end of the bibliographic entry. This is working now. You are able to switch this feature on and off by using the annote option in the preamble or in your local jurabib.cfg file.

textedition = In v0.51e this field was removed again, because enclosing the edition in curly v0.51e CHANGED ! brackets will have the same result.

6.2.6 Citing Juridical Dissertations and the Like

dissyear = Of course one can cite juridical dissertations as normal dissertations, but a juridical dissertation may have been published as a book as well. In the first case, the entry

type @JURTHESIS (or @PHDTHESIS/@MASTERSTHESIS) should be used. In the second

case, using the entry type @BOOK should be preferred. For this reason, a new field dissyear has been created, which enables one to specify the year in which a book appeared as doctoral dissertation. Furthermore, if dissyear is present, fields school and type are available for @BOOK as well. If dissyear is not present, type and school are ignored. For example:

  @BOOK{alexy,
    author      = {Alexy, Robert},
    title       = {Theorie der Grundrechte},
    year        = 1985,
    address     = {Baden-Baden},
    type        = {Habil.},
    school      = {G{\"o}ttingen},
    dissyear    = 1984
  }

shows the following:

Alexy, Robert: Theorie der Grundrechte, Baden-Baden 1985 (also Habil. Göttingen 1984)

Suppose this work had not been published as a book, the following entry would have been appropriate::

  @PHDTHESIS{alexy,
    author      = {Alexy, Robert},
    title       = {Theorie der Grundrechte},
    year        = 1984,
    type        = {Habil.},
    school      = {G{\"o}ttingen}
  }

and would have shown the following:

Alexy, Robert: Theorie der Grundrechte, Habil. Göttingen 1984

The default for this type is “Jur. Diss.”:

  @PHDTHESIS{alexy,
    author      = {Alexy, Robert},
    title       = {Theorie der Grundrechte},
    year        = 1984,
    school      = {G{\"o}ttingen}
  }

Alexy, Robert: Theorie der Grundrechte, Jur. Diss. Göttingen 1984

\SS The command \SS redefines the standard L^{A}TEX-command in such a way that it now produces two section marks with reduced space in between the two. Compare \S\S: §§ and \SS: §§

7 Remaining things …

Some helpful commands, which were introduced into the package without any announcement.

pages=format (Primarily for the humanities) You can switch on preformatting of the pages given by the pages-field. You are able to write pages = {22-34} instead of pages = {pp.~22-34}. Not enough, you are able to do the same with the pages given by the optional argument of your \cite command. If you want to give something else as page(ranges)s or if you want to add something after a page(range), please use the \nopage and \pageadd command (see below for an example). jurabib will insert the appropriate and—if you are using babel—it will use the chosen main document language. jurabib makes a difference between a page and a pagerange. If you like to change the defaults:

     \AddTo\bibsenglish{%
        \def\jbpagename{page}%
        \def\jbpagesname{pages}%
     }

0.51g NEW ! With version 0.51g you can use separate macros for the bibliography. The macros are named \bibpagename and \bibpagesname. Please note that the definitions of \bibpage[s]name are the same as for \jbpage[s]name. This means, if you are redefining \jbpage[s]name, this will have effect on the

meaning of \bibpage[s]name unless you are redefining \bibpage[s]name

separately.

Let’s take a look what jurabib can do to minimize your work^{7}:

  \cite[45]{<key>}  . . . , p. 45
  \cite[45--47]{<key>}  . . . , pp. 45–47
  \cite[45, 47 and 49]{<key>}  . . . , pp. 45, 47 and 49
  \cite[45f]{<key>}  . . . , pp. 45f.
  \cite[45ff]{<key>}  . . . , pp. 45ff.
  \cite[\nopage{I, III and IV}]{<key>}  . . . , I, III and IV
  \cite[13,\pageadd{something text}]{<key>} . . . , p. 13, something text

          ^No whitespace here!!!

pages=test By default, page(range)s, which are given via the pages-field in the .bib-file, are suppressed in the citation. With pages=test it will be tested, if a page(range) is given by the optional argument of the \cite command. If so, that one will be used. If there’s no optional page(range) given, the one from the .bib-file will be used.

pages=always The page(range) given by the .bib-file are printed always.

0.51c NEW ! hypercite=false Disables the automated conversion of citations into hyperlinks when using the hyperref package.

\jbdotafterbibentry places a dot at the end of each entry in the bibliography.

0.5f NEW ! \jbdotafterendnote places a dot at the end of each endnote (if you are using endnotes.sty).

\jbedseplikecite Separation of the editors in the bibliography will be the same like in the citation.

\jbdisablecitationcrossref Disables crossrefs, which aren’t @INCOLLECTIONs.

\jbsilent Suppresses all jurabib warnings.

\formatpages formerly known as \formatarticlepages This command al-

lows you to determine the appearance of page(range)s of citations of all types. jurabib is able to extract the starting page from a given pagerange (in the database). Since v0.5f this command can be used with all publication types. You can specify a list of publication types by using the first manda-

0.51o CHANGED ! tory argument. This command takes two ! optional and three mandatory arguments:

\formatpages[⟨after start page separator⟩][⟨before start page separator⟩]{⟨typelist⟩}{⟨left delim⟩}{⟨right delim⟩}

If you type \formatpages[: ]{article,periodical}{(}{)}, jurabib will

format the citation \cite[48]{⟨key⟩} with the follwowing database entry:

    @ARTICLE/PERIDOCAL{broxja,
      author      = {Hans Brox},
      title       = {Die Anfechtung bei der Stellvertretung},
      journal     = {JA},
      language    = {german},
      year        = 1980,
      pages       = {45--60},
      address     = {M{\"u}nchen}
    }

like that:

Brox, JA 1980, 45: (48).

If you want format only the pages given by the optional argument of \cite, you can leave the optional arguments empty:

\formatarticlepages{article}{[}{]}

Brox, JA 1980, [48].

0.52b NEW ! If you are using both \formatpages and pages=format, it is default now to suppress formatting of pages given by the optional argument of \cite, e.g. you have written \formatpages[, ]{article}{}{} into your preamble and you are using pages=format too, then the ouput will look like:

Brox, JA 1980 p. 45, 48.

7This rewritten option was inspired by pageranges.sty, available at CTAN, and completely contributed by Stefan Ulrich. Thanks a lot.

If you want to have a formatted second page number, you can use

     \jbnoformatafterstartpagefalse in the preamble:

Brox, JA 1980 p. 45, p. 48.

0.52h NEW ! \jbfirstcitepageranges If you have given a pagerange via the pages-field for @ARTICLE or @PERIODICAL type, this pagerange will be printed out for first (using citefull=first) and full citations done by using \[foot]fullcite. This works independently from the pages options ! If you add a page using the optional argument of the \[foot]cite command, this page will be added after the pagerange, separated by “here:”, which is represented by the macro

     \herename:

[…], p. 45, here: p. 48.

Redefinition as usual:

         \AddTo\bibsenglish{%
             \def\herename{there:}%
         }

8 The Configuration file jurabib.cfg

That file can be used to save redefinitions and options. The name has to be jurabib.cfg and it has to live in the working directory or in the same directory where jurabib lives.

   \jurabibsetup{%
     authorformat=smallcaps,
     commabeforerest,
     titleformat=colonsep,
     bibformat={tabular,ibidem,numbered}
   }

9 Options for Other Academic Disciplines

The following options are not intended primarily for juridical work, but satisfy the needs of, among others, historians, philosophers, etc. They are the (provisional) answers to the most different questions which were proposed to me. I hope them to be useful. Suggestions for improvement are welcome.

ibidem=name By this option, more extensive data will be inserted then by using ibidem/ibidem=strict. If this option is active, the complete name of the author will be given (If not the option authorformat=reducedifibidem is active, because then, only the last name will be given). This option is intended for use together with citefull=first, and therefor, that option will be activated automatically. If an author is cited with multiple works, it may happen that jurabib automatically switches to the next option, for guaranteeing the unambiguity of the citation. You will find a hint to this in the .log-file. ibidem=name&title Just like ibidem=name, but in this case, the title will be given

as well. Also here, the option citefull=first will be active.

This is a survey to the options explained above:

\footcite..       ibidem=name                        ibidem=name&title

..{brox:bgb}

1

Brox, Hans: Allgemeiner Teil des Bürgerlichen Gesetzbuches.

20. Auflage, Köln, Berlin, Bonn, München 1996.

1

Brox, Hans: Allgemeiner Teil des Bürgerlichen Gesetzbuches.

20. Auflage, Köln, Berlin, Bonn, München 1996.

..{brox:bgb}   ^{2}Brox, Hans, ibid.  ^{2}Brox, Hans: BGB AT, ibid.
..{brox:bgb}   ^{3}Brox, Hans, ibid.  ^{3}Brox, Hans: BGB AT, ibid.

..{oellers}

4

Oellers, Bernd: Doppelwirkung im Recht ? AcP 169 [1969].

4

Oellers, Bernd: Doppelwirkung im Recht ? AcP 169 [1969].

..{brox:bgb}   ^{5}Brox, Hans, ibid.  ^{5}Brox, Hans: BGB AT, ibid.
..{brox:bgb}   ^{6}Brox, Hans, ibid.  ^{6}Brox, Hans: BGB AT, ibid.

Pagebreak

..{brox:bgb}   ^{7}Brox, Hans, ibid.  ^{7}Brox, Hans: BGB AT, ibid.
..{brox:bgb}   ^{8}Brox, Hans, ibid.  ^{8}Brox, Hans: BGB AT, ibid.

now with authorformat=citationreversed:

..{brox:bgb}   ^{9}Hans Brox, ibid.  ^{9}Hans Brox: BGB AT, ibid.

or with authorformat=reducedifibidem:

..{brox:bgb}   ^{10}Brox, ibid.  ^{10}Brox: BGB AT, ibid.

0.5f NEW ! ibidem=name&title&auto This option can be useful for often repeated citations of different work of the same author. For the first citation the full entry is printed (citefull=first is automatically used). As long as the same work from an author is cited, only the name of the author will be used (this is

equal to ibidem=name).

If the work is cited again some footnotes later, name and title will be printed out (ibidem=name&title). This will avoid ambiguity if an author is cited with more than one work. If the immediately following citations are from the same author, only the name will be printed out.

If you are using ibidem=name only, jurabib will check if the citations seems to be ambiguous and then, ibidem=name&title&auto will be used automatically. You will find a remark in the .log file.

Here a survey:

 \footcite..       ibidem=name&title&auto

 ..{brox:bgb}

1

Brox: Allgemeiner Teil des Bürgerlichen Gesetzbuches.

20. Auflage, Köln, Berlin, Bonn, München 1996.

..{brox:schr}   ^{2}Brox: Besonderes Schuldrecht. 20. Auflage, München 1995.

 ..{brox:bgb}    ^{5}Brox, ibid.

now another work is cited !

..{brox:schr}   ^{6}Brox: SchR BT, ibid.

 ..{brox:schr}   ^{7}Brox, ibid.

… and we are switching back again !

..{brox:bgb}    ^{8}Brox: BGB AT, ibid.

 ..{brox:bgb}    ^{8}Brox, ibid.

Changing the sequence of “ibid.” is possible by redefining the commands

\ibidemname and \ibidemmidname:
      \AddTo\bibsenglish{%
         \renewcommand{\ibidemname}{Ibid.}
         \renewcommand{\ibidemmidname}{ibid.}
      }
\ibidemname is used by the options ibidem=strict, ibidem=strictdoublepage

and ibidem=nostrict and appears at the beginning of the—suppressed—citation (It is therefore possible to capitalize it). \ibidemmidname however, appears together with the options ibidem=name and ibidem=name&title and can be written in lowercase—which depends on your preference.

0.51 NEW ! Now it’s possible to make a difference between a subsequent citation with same page(s) and a subsequent citation with different page(s). For the first case, the

macros \samepageibidemname or \samepageibidemmidname will be used inter-

nally. Its definiton is the same as for \ibidemname or \ibidemmidname by default.

For the second case, the macros \diffpageibidemname or \diffpageibidemmidname

are used. If you like to make a difference between these two cases, you can redefine the macros to your needs. I don’t know useful abbreviations for the English language, so I will demonstrate it with a dummy:

     \renewcommand*{\samepageibidemname}{[same pages]}

A short table will explain more:

               \footcite[45]{broxbgb}  ⟶ ^{1}Brox, 45.
               \footcite[45--47]{broxbgb} ⟶ ^{2}Ibid., 45–47.
               \footcite[45--47]{broxbgb} ⟶ ^{3}[same pages]
               \footcite[45f]{broxbgb}  ⟶ ^{4}Ibid., 45 f.

citefull=first By aid of this option, it is possible to show the full bibliography entry in the first citation. For all subsequent citations, a short form will be used. To force a full citation in later citations, use the commands \fullcite and \footfullcite. With this option, editors appear before author’s names, separated by ‘in”. In order to get uniform citation and to avoid confusion of the reader, this option automatically activates the option cofirstsep=in and thus activates options which are presupposed by that option. Apart from that, the howcited-options are deactivated. citefull=chapter switches on citefull=first automatically and resets each

chapter. citefull=section switches on citefull=first automatically and resets each

section. citefull=all This option shows all citations as full citations. It switches also

all separators to cofirstsep=in. The howcited-options are deactivated. However, combination with ibidem is possible.

see Because those outside law studies don’t need the second optional argument of the \cite* commands, by this option, one can add phrases like “See” or “Compare” before the citation. This option works globally.

natoptargorder Hereby, the sequence of the optional parameters is reversed, for instance because of making the document compatible with natbib.sty (before writing it).

crossref=dynamic You can use the other crossref-options together with that option to achieve crossrefs, which are different in their length—longer, if a work is cited the first time—shorter if it is cited again. Please compare the lonely use of crossref=dynamic (for better understanding you may find the crossrefs inside square brackets):

1 Lincoll, Daniel D.: Semigroups of Recurrences. In [Lipcoll/Lawrie/Sameh: High Speed Computer and Algorithm Organization]. 2 Lincoll, Daniel D.: Semigroups of Recurrences. In [Lipcoll/Lawrie/Sameh].

with the combination of crossref=dynamic with crossref=long:

1 Lincoll, Daniel D.: Semigroups of Recurrences. In [Lipcoll, David J./Lawrie, D. H./Sameh, A. H. (eds.): High Speed Computer and Algorithm Organization. 3rd edition, New York: Academic Press, September 1977 (Fast Computers 23)]. 2 Lincoll, Daniel D.: Semigroups of Recurrences. In [Lipcoll/Lawrie/Sameh: High Speed Computer and Algorithm Organization].

crossref=normal (Default) Hereby, cross references specified by the special field crossref in the bibliography are displayed with author (shortauthor prevailing over author) and title (shorttitle if available, else title).

crossref=short If no ambiguities result, title is left out with this option. In case

a title is needed, shorttitle prevails over title. crossref=long With this option, the cross reference will be displayed as full

citation. lookat Hereby, references to footnotes are enabled which contain the full cita-

tion to which is referred. This is possible only when using the \footcite command in connection with the option citefull=first. This may be useful for articles which do not contain bibliographies. For this purpose, the command

\nobibliography            \nobibliography{⟨bibfile⟩}

is available, which suppresses bibliography generation.^{8} In later citations, short forms of citation are used, accompanied by a hint to the footnote in which the full citation can be found.^{9} Please notice that for correct parsing of the references with lookat, after running BibTEX, three L^{A}TEX-runs are necessary!

8This command is primarily for use with lookat, but it doesn’t require lookat. But one of the citefull-options has to be enabled. Thanks to Stefan Ulrich.

9Example: Brox: BGB AT (wie Anm. ⟨Nr.⟩), Rn. 168.

In case you want to use the package varioref or fancyref, then instead of \ref the command \vref will be used, resulting in different references, in particular if the full citation is one or two pages away. In case you want to use the package varioref or fancyref in your document, without consequences

\jbignorevarioref for your citations, you can put the command\jbignorevarioref in the preamble of your document.

It is possible to customize the behavior of the lookat option by the following commands:

   \renewcommand{\jblookatprefix}{(see footnote~}
   \renewcommand{\jblookatsuffix}{)}

lookat may be used in connection with the ibidem-options, but such is not recommended.

When using \cite commands in the main text (outside footnotes!) and

CAVE ! \footcite commands (or \cite commands inside footnotes) errors may occur from alphanum or—if used—varioref. alphanum may complain in the following way:

    ! Package alphanum Error: Self-reference detected!.
       [...]
    ?

In this case, you have cited a work in the main text for the first time, to which one refers in the same section by a \footcite command. Because the first citation is not inside a footnote, the section number will be taken as reference label. Because by using the option lookat, following citations refer to the first citation, and this citation is in the same section, alphanum complains as mentioned above.

Another error report which is caused by the same problem might look as follows:

    ! Extra }, or forgotten \endgroup.
    \J@refP ...nta #1\,\J@INumberRoot {#1}{#2}
       [...]
    ?

The solution is to put at least the first citation inside a footnote (whether one uses \footcite or \cite inside a \footnote makes no difference.).

The lookat-option cannot be used with documents based on the book- or report-class or their derivatives. Use of this option with the footnpag.stypackage is equally impossible.

Using lookat in connection with book- or report-classes is possible by the remreset-package. This package disables resetting the footnote counter at the start of a new chapter, thus enabling unambiguous references. In order to do so, you should insert the following in your preamble:

    \usepackage{remreset}
    \makeatletter
    \@removefromreset{footnote}{chapter}
    \makeatother

In oder to gain consequent results, you should also properly set the counters of images and tables:

    \usepackage{remreset}
    \makeatletter
    \@removefromreset{footnote}{chapter}
    \@removefromreset{figure}{chapter}
    \renewcommand{\thefigure}{\@arabic\c@figure}
    \@removefromreset{table}{chapter}
    \renewcommand{\thetable}{\@arabic\c@table}
    \makeatother

idem This new option is very similar to the ibidem option, but there is not the whole citation replaced by an abbreviation. If the author is cited again, his name will be replaced by “Idem” or “idem”. idem is working together with all ibidem options. The following values are possible: idem (same as

idem=strict), idem=strictdoublepage and idem=nostrict. The behav-

ior of idem at page breaks is the same as for the ibidem options, therefore I did not make any survey. Redefinitions as usual:

       \AddTo\bibsenglish{%
          \renewcommand*{\idemname}{Eadem}
          \renewcommand*{\idemmidname}{eadem}
       }

\noidem \noidem is working analogous to \noibidem and disables the idem mechanism for the following citation.

opcit This option is at experimental stage and places the abbreviation “op. cit.” (opere citato: already cited) in the citation. An example:

1 Aamport, Gnats and Gnus (1986), p. 25. […] 5 Aamport, op. cit., p. 37.

If you like to modify:

      \renewcommand*{\opcit}{\textit{op.\,cit.}}

opcit=chapter opcit can be resetted at the begin of each chapter/section with the

opcit=section  values chapter/section. This is analogous to citefull=chapter or
citefull=section.

10 jurabib-specific fields

jurabib.bst:

        url
        urldate
        volumetitle
        dissyear
        howcited
        juratitle
        juraauthor
        shorttitle
        shortauthor
        ssedition
        titleaddon
        booktitleaddon
        editortype
        sortkey
        annote
        gender
        translator
        flanguage
        oyear
        oaddress
        opublisher
        updated
        totalpages
        founder

11 Linguistic Stuff

Because of increasing use of the jurabib package by people outside law studies, I have reworked the BibTEX-styles in such a way that it is possible to switch between different languages.

language = It is now possible, to specify the language of a certain BibTEX-entry by providing the field language with the appropriate language. At the moment, only English, German, French, Dutch, Spanish and Italian are implemented:

      @INCOLLECTION{incollection-crossref,
          author      = {Daniel D. Lincoll},
          title       = {Semigroups of Recurrences},
          pages       = {179--183},
          language    = {english}
      }

If language = {⟨other language ⟩} is specified here, then the hyphenation patterns for that entry will be modified, not the keywords like “editor” and so on ! For the correct translation of the keywords jurabib will detect the used main language (e.g. if you are using the babel-package or one of the following packages: german, french, frenchle, pmfrench, mlp).

jurabib doesn’t switch the keywords, but the hyphenation patterns for each bibliographic entry, for which a language is given !

You should notice, that (for German users: contrary to the bibgerm-package) only the relevant entries (which differs from the main language) should be provided with language-fields. You can modify the defaults by using the \AddTo functionality:

    \AddTo\bibsgerman{\def\editorname{ed.}}

The macro for English is \bibsenglish, for French \bibsfrench, for Dutch

\bibsdutch, for Spanish \bibsspanish.

If I’ve forgotten something which has to do with language support, please e-mail me. The same goes for supporting more languages.

12 Across the bounderies

12.1 jura.cls

As I said already, use of jura.cls is possible.

12.2 bibtopic.sty

jurabib.sty is already for a long time—under certain conditions—compatible with bibtopic.sty: All short titles should be specified, because otherwise, ambiguous citations may occur. At this moment, it is not possible to use automatic generation of short titles when separated bibliography files are in use. Therefore, the option titleformat=all is activated automatically, if bibtopic.sty has been loaded. Note: You should use at least version 1.0j.

12.3 hyperref.sty

jurabib is compatible with hyperref.^{10} However, incompatibility still exists with alphanum, of which the result is that the \ref mechanism of alphanum functions only to a limited extent. That means primarily that the command \ref* provides relative references only. If you want to use alphanum and hyperref without jurabib, you should put the following in your preamble:

   \makeatletter
     \let\J@SetCurrent\relax
     \def\toclevel@lvla{0}\def\toclevel@lvlb{1}
     \def\toclevel@lvlc{2}\def\toclevel@lvld{3}
     \def\toclevel@lvle{4}\def\toclevel@lvlf{5}
     \def\toclevel@lvlg{6}\def\toclevel@lvlh{7}
     \def\toclevel@lvli{8}\def\toclevel@lvlj{9}
     \def\toclevel@lvlj{10}\def\toclevel@lvll{11}
     \newcommand*{\theHlvla}{\J@Number}\newcommand*{\theHlvlb}{\J@Number}
     \newcommand*{\theHlvlc}{\J@Number}\newcommand*{\theHlvld}{\J@Number}
     \newcommand*{\theHlvle}{\J@Number}\newcommand*{\theHlvlf}{\J@Number}
     \newcommand*{\theHlvlg}{\J@Number}\newcommand*{\theHlvlh}{\J@Number}
     \newcommand*{\theHlvli}{\J@Number}\newcommand*{\theHlvlj}{\J@Number}
     \newcommand*{\theHlvlk}{\J@Number}\newcommand*{\theHlvll}{\J@Number}
     \renewcommand{\J@LongToc}[2][]{
       \@startsection{lvl\alph{tiefe}}{\number\value{tiefe}}{0pt}
       {\ifnum\value{tiefe}=1 -4ex plus-1,5ex minus-0,ex\else
       -2,7ex plus-0,8ex minus-0,2ex\fi}{\ifnum\value{tiefe}>7

10Stefan Ulrich was behind this feature as well.

       -1em plus-0,5em\relax\else 0,6ex plus0,3ex minus0,1ex\fi}
       {\sectfont\csname lvl\alph{tiefe}style\endcsname}[#1]{#2}
     }
   \makeatother

12.4 babel.sty

jurabib is compatible with babel.

12.5 chapterbib.sty

jurabib is compatible with chapterbib.

12.6 bibunits.sty

jurabib is fully compatible with bibunits, you should use v2.1n or higher.

12.7 index.sty

If you are using the french-, pmfrench- or the frenchle-packge, you are not able to use this feature at the moment.

With the option authorformat=indexed you can index all cited authors. If you want to generate a separate author index, it’s possible to use the index package by David M. Jones, which is part of the camel bundle.

   [...]
  \usepackage{index}
  \newindex{default}{idx}{ind}{Index}    % for the normal Index
  \newindex{aut}{adx}{and}{Authorindex}  % for the new author Index
  \renewcommand{\jbindextype}{aut}       %
   [...]
  \begin{document}
   [...]
  \printindex                            % for the normal Index
  \printindex[aut]                       % for the new author Index
  \end{document}

Please note that \jbindextype has to contain the same value as the first argument

of \newindex.

To generate the index, run:

   makeindex -o datei.and datei.adx

Then you have to run L^{A}TEX again. For further explanations please take a look at the index package documentation.

12.8 endnotes.sty

0.51 NEW ! jurabib is now compatible to endnotes. With a simple \usepackage{endnotes} in your preamble all your \foot[full]cite citations are converted into endnotes, which are appearing where you typed \theendnotes. Citations, which are enclosed in \footnote commands or normal \cite commands, were not converted by default ! If you need this feature, please use the option ‘citetoend=true’ or consult the documentation of the endnotes package.

Endnotes does not have a closing dot by default. If you would have a closing period, please use \jbdotafterendnote in the preamble of your document.

13 Cooperators

This package would have never been come into existence without the powerful support of Stefan Ulrich, Andreas Stefanski and Oren Patashnik. It is especially due to Stefan Ulrich that the package now exists in its present form and not stumbled in its provisional stage of development. To him, my special thanks. Heiko Oberdiek provided some valuable hints. Andreas Stefanski was my unexhaustible tester and advisor in juridical formalities. Not to omit Oren Patashnik, who provided important parts of the BibTEX-styles which enable the dynamic generation of the juridical shorttitles. Bernard Gaulle has done a lot for compatiblity with the french packages and gave many other helpful hints, especially about the linguistic stuff. I want to thank Maarten Wisse for translating the documentation, for his patience and for a lot of helpful hints during implementation of the basic humanities features. Peter Flynn and Pádraig de Brún also suggested a lot of new features for the humanities. And there are lots of people who worked as beta testers and reported bugs: Alexander Wisspeintner, Andreas K. Foerster, Arne Engels, Axel Sodtalbers, Bastian Kruse, Christian Folini, Christian Meyn, David Feest, Daniel M. Grisworld, Hélène Fernandez, Henning Eiden, Holger Pollmann, Hubert Selhofer, Ivan Blatter, Jean-Pierre Drucbert, Joachim Trinkwitz, Max Dornseif, Moritz Moeller-Herrmann, Nikolai Warneke, Olaf Meltzer, Oliver Schilling, Peter Wuesten, Ralph Sinkus, Rebekka Rieger, Robert Goulding, Thorsten Manegold and Tilman Finke.

14 Response requested …

Questions, recommendations and critique or the like can be sent to: jb@jurabib.org